OpenStudy (anonymous):

The concentration C(t) of a certain drug in the blood stream after t minutes is given by the formula C(t) =0.09(1-e^-0.2(t)). What is the concentration after 18 minutes? I don't understand the "e".... please help.

OpenStudy (anonymous):

\[\large C(t)=0.09\left(1-e^{-0.2t}\right)\] \(e\) is a mathematical constant, roughly equal to 2.718. If you have a graphing calculator, you should have a button designated for it. To find the concentration after 18 minutes, you plug in \(t=18\): \[\large C(18)=0.09\left(1-e^{-0.2\times18}\right)=\cdots\]

OpenStudy (anonymous):

Thank You!!

OpenStudy (anonymous):

yw